# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

### v1

@@ 8.3 Distortion Identification Analysis
Table 6: The total and proportion of occurrences and distortions for body parts in the $400$ videos generated by each model. Red represents the minimum value in a column.
@@
Face Arms Torso Legs Feet None / Total
Latte [55] 358 / 250 324 / 310 329 / 140 89 / 56 48 / 38 21 / 794
89.5 / 69.8 81.0 / 95.7 82.3 / 42.6 22.3 / 62.9 12.0 / 79.2 5.2 / 69.2
AnimateDiff [22] 373 / 245 352 / 316 357 / 88 183 / 67 108 / 63 22 / 779
93.2 / 65.7 88.0 / 89.8 89.2 / 24.6 45.8 / 36.6 27.0 / 58.3 5.5 / 56.7
Gen-2 [16] 383 / 250 357 / 329 361 / 134 138 / 86 93 / 72 10 / 871
95.8 / 65.3 89.2 / 92.2 90.2 / 37.1 34.5 / 62.3 23.3 / 77.4 2.5 / 65.4
StreamingT2V [26] 361 / 304 366 / 342 348 / 199 170 / 127 105 / 92 15 / 1064
90.2 / 84.2 91.5 / 93.4 87.0 / 57.2 42.5 / 74.7 26.3 / 87.6 3.7 / 78.8
VideoCrafter2 [9] 382 / 324 350 / 295 365 / 117 98 / 60 64 / 49 11 / 845
95.5 / 84.8 87.5 / 84.3 91.2 / 32.1 24.5 / 61.2 16.0 / 76.6 2.8 / 67.1
Open-sora-plan [40] 379 / 333 376 / 371 367 / 214 134 / 84 68 / 60 9 / 1062
94.8 / 87.9 94.0 / 98.7 91.7 / 58.3 33.5 / 62.7 17.0 / 88.2 2.2 / 80.2
MagicTime [98] 320 / 292 309 / 302 303 / 177 248 / 177 216 / 187 66 / 1135
80.0 / 91.2 77.2 / 97.7 75.7 / 58.4 62.0 / 71.4 54.0 / 86.6 16.5/ 81.3
StableVideo [1] 393 / 289 378 / 323 381 / 127 128 / 60 72 / 47 2 / 846
98.3 / 73.5 94.5 / 85.4 95.3 / 33.3 32.0 / 46.9 18.0 / 65.3 0.5 / 62.6
We calculate the proportion of occurrences and distortions for each body part of AGVs for different T2V models, as detailed in Table 6. For body occurrences, StableVideo performs the best, with the highest proportion of main body parts (face, arms, torso) present among these T2V models, and only $0.5\%$ of videos generated by StableVideo lack any human depiction. In contrast, MagicTime performs the worst, with $16.5\%$ of videos missing any human body parts. In terms of distortions, AnimateDiff and StableVideo exhibit better performance, with VideoCrafter2 and Gen-2 also showing comparable results. In contrast, Open-Sora and MagicTime show significant deficiencies, with distortions in certain body parts accounting for over $95$%. These findings closely align with the subjective quality scores of corresponding T2V models.
Moreover, an observation of significant concern is that $16.5\%$ of videos generated by MagicTime lack any depiction of human body parts, indicating a shortfall in the model’s ability to maintain consistency between textual descriptions and visual content. StableVideo demonstrates the most robust performance in this regard.
@@
8.4 The Relationship between Semantic Distortion Identification and Visual Quality Scoring
@@
Figure 13: The correlation between the semantic distortion identification and subjective quality.
@@
8.4.1 Body Parts Distortion
Overall, the quality of T2V models across the three evaluated dimensions demonstrates a certain degree of correlation with the distortions in the five body parts. Among these, facial distortion has the most significant impact on human appearance quality, followed by distortions in the torso and legs. This pattern aligns with the natural human perception process, which prioritizes facial recognition before focusing on the torso and other body parts.
Of the three quality dimensions, HA Quality exhibits the strongest correlation with body part distortions, followed by AC Quality, with OV Quality showing the weakest correlation. This suggests that body part distortions most directly influence perceptions of human appearance quality, with a lesser effect on the continuity of human actions, and the least impact on overall video quality.
Interestingly, for certain models, such as VideoCrafter2, quality scores show a negative correlation with distortions in the arms and feet. This implies that viewers’ attention is drawn more to prominent areas like the face and torso, making distortions in less visually critical body parts less influential.
8.4.2 Body Parts Presence
@@
Interestingly, the presence of legs and feet often correlates negatively with all three quality dimensions, suggesting that their inclusion tends to degrade overall video quality. This underscores the current limitations of T2V models in accurately generating human legs and feet, as distortions in these areas significantly impact viewers’ subjective perception of video quality across all dimensions.
9 Experiment Settings
9.1 Evaluation Metrics
@@
$PLCC=\frac{\sum_{n=1}^{N}(y_{n}-\bar{y})(\hat{y}_{n}-\bar{\hat{y}})}{\sqrt{\sum_{n=1}^{N}(y_{n}-\bar{y})^{2}\sum_{n=1}^{N}(\hat{y}_{n}-\bar{\hat{y}})^{2}}},$ (13)
@@
9.2 Compared Quality Metrics
9.2.1 Image Quality Assessment Methods
@@
NIQE [58] is a no-reference image quality assessment (IQA) model that evaluates the quality of images based on natural scene statistics (NSS) without relying on human-rated training data. For the Human-AGVQA dataset, we applied NIQE to assess image quality using selected video frames.
@@
BRISQUE [57] leverages directional contrast features extracted from the Curvelet domain to assess the quality of contrast-distorted images. For the Human-AGVQA dataset, we utilized this approach to evaluate image quality based on its directional contrast features.
@@
HyperIQA [68] introduces a self-adaptive hyper network for blind image quality assessment (BIQA) in the wild, addressing diverse content and distortion types in real-world images. For the Human-AGVQA dataset, we applied HyperIQA to assess image quality by adaptively learning perception rules and predicting quality scores.
@@
UNIQUE [101] is a unified uncertainty-aware BIQA model designed to handle synthetic and realistic distortions. The model predicts both quality scores and associated uncertainties, enabling a probabilistic interpretation of image quality. For the Human-AGVQA dataset, we used UNIQUE to assess video frame quality based on its learning-to-rank framework.
@@
MUSIQ [34] introduces a multi-scale image quality Transformer designed to handle full-resolution images with varying sizes and aspect ratios. The model uses hash-based 2D spatial embedding and scale embedding to capture positional and scale information effectively. For the Human-AGVQA dataset, we applied MUSIQ to evaluate video frame quality by leveraging its multi-scale feature extraction capability.
@@
StairIQA [72] introduces a staircase network that integrates features from multiple convolutional layers of a CNN to capture better both low-level visual details and high-level semantic information for in-the-wild image quality assessment. For the Human-AGVQA dataset, we used the official implementation and adopted this approach to evaluate image quality without modifications.
@@
CLIP-IQA [81] is a visual perception assessment model that utilizes the pre-trained CLIP framework to evaluate both the quality and abstract attributes of images. It introduces an antonym prompt pairing strategy (e.g., Good photo vs. Bad photo) to mitigate linguistic ambiguity and employs cosine similarity to predict image quality scores. For the Human-AGVQA dataset, we used the official implementation, applying CLIP-IQA to assess video frame quality.
@@
LIQE [102] is a multitask learning framework for blind image quality assessment (BIQA) that leverages vision-language correspondence. LIQE uses CLIP for feature embeddings and optimizes tasks jointly by marginalizing probabilities over a textual template that describes scene, distortion, and quality. For the Human-AGVQA dataset, we used the official implementation, applying the quality assessment textual template to evaluate the video frame quality.
@@
MA-AGIQA [84] introduces a framework for AI-generated image quality assessment by combining large multi-modality models (LMMs) with traditional deep neural networks (DNNs). For the Human-AGVQA dataset, we used the official implementation, applying MA-AGIQA to evaluate the video frame quality without modifications.
9.2.2 Video Quality Assessment Methods
@@
TLVQM [38] is a two-level video quality model that first computes low-complexity features across the entire video and then selects representative frames based on these features to extract high-complexity features.
@@
RAPIQUE [78] is a video quality assessment model that combines spatial-temporal scene statistics with high-level features extracted using deep CNNs. For the Human-AGVQA dataset, we used the official implementation, applying sparse frame sampling and integrating spatial and temporal features to evaluate video quality.
@@
VIDEVAL [77] applies a feature selection strategy based on efficient blind VQA models. Using the official open-source code, we converted Human-AGVQA videos from RGB to YUV420 format for feature extraction.
@@
Patch-VQ [96] is a video quality model that extracts spatial and temporal features from video frames and patches to analyze the relationship between local distortions and overall video quality. For the Human-AGVQA dataset, we used Patch-VQ’s official implementation, dividing each video into space-time patches for feature extraction and quality prediction.
@@
SimpleVOA [71] uses an end-to-end network to extract spatial features directly from video frames and motion features to capture temporal distortions. A pre-trained SlowFast model processes 8 uniformly sampled frames rescaled to a height of 520. We used the official SimpleVOA model and finetuned it on Human-AGVQA.
@@
FAST-VQA [87] introduces a grid mini-patch sampling (GMS) strategy to balance local quality assessment with global context while reducing computational costs for high-resolution videos. We used the official FAST-VQA-B model and finetuned it on Human-AGVQA.
@@
DOVER [88] is a video quality evaluator that assesses quality from technical and aesthetic perspectives. We used the official code without modifications.
@@
T2VQA [39] is a transformer-based model designed to evaluate text-to-video quality, focusing on text-video alignment and video fidelity. It uses BLIP and Swin-T for feature extraction, followed by feature fusion through a cross-attention module and quality regression using a large language model. On the Human-AGVQA dataset, we used the official T2VQA model and finetuned it on Human-AGVQA.
Table 7: Details of the model structure for the proposed model.
@@
UGVQ [103] is a unified framework for video quality assessment, focusing on spatial quality, temporal coherence, and text-to-video alignment. It utilizes features from CLIP for text and visual representations and SlowFast for motion representation. For the Human-AGVQA dataset, we employed UGVQ’s feature extraction and fusion modules to evaluate video quality comprehensively across these three dimensions.
EvalCrafter [54] is an evaluation framework for text-to-video (T2V) generative models, focusing on video quality, text-video alignment, motion quality, and temporal consistency. It uses a benchmark of 700 prompts and 17 evaluation metrics to analyze T2V model performance. For overall video quality, we directly used implementation code in EvalCrafter on the Human-AGVQA dataset without specific changes.
Q-Align [89] introduces a methodology for teaching large multi-modality models (LMMs) to assess video quality using discrete text-defined levels such as excellent, good, and poor. It converts mean opinion scores (MOS) into these rating levels for training and infers scores by weighting the predicted probabilities of the levels. For the Human-AGVQA dataset, we used the official code for inference without modifications.
9.2.3 Action Quality Assessment Methods
USDL [74] is an uncertainty-aware score distribution learning approach for action quality assessment (AQA), treating actions as instances associated with score distributions. For the Human-AGVQA dataset, since all videos contain more than 16 frames, each video is evenly divided into ten segments. Feature extraction is performed using an I3D backbone pre-trained on Kinetics. The final scores are normalized linearly, scaling the raw scores based on their minimum and maximum values to a range of [0, 100]. These normalized scores are then used to construct Gaussian distributions with the normalized value as the mean.
ACTION-NET [99] is a hybrid attention network tailored for action quality assessment (AQA) in long videos. It combines dynamic video information with static postures of action subjects in selected frames. For the dynamic stream, frames are sampled at a rate of 4 frames per second. For the static stream, the first, middle, and last frames are sampled, and the action subjects are cropped using the original detection algorithm.
Table 8: Details of the hyper-parameters for the model training.

### v2

@@ B.3. Distortion Identification Analysis
Table 6. The total and proportion of occurrences and distortions for body parts in the $400$ videos generated by each model. Red, blue, and green represent the first, the second, and the third minimum values in the column, respectively.
@@
Face Hands Arms Torso Legs Feet Total None
Latte (Ma et al., 2024) 358 / 284 298 / 295 292 / 166 329 / 174 89 / 65 48 / 41 1414 / 1025 21
89.5 / 69.8 74.5 / 95.7 73.0 / 95.7 82.3 / 42.6 22.3 / 62.9 12.0 / 79.2 - / 72.5 5.25
AnimateDiff (Guo et al., 2024) 373 / 272 312 / 299 340 / 116 357 / 124 183 / 82 108 / 67 1673 / 960 21
93.3 / 65.7 78.0 / 89.8 85.0 / 89.8 89.2 / 24.6 45.8 / 36.6 27.0 / 58.3 - / 57.4 5.25
Gen-2 (Esser et al., 2023) 383 / 269 328 / 317 362 / 173 361 / 176 138 / 109 93 / 88 1665 / 1132 10
95.8 / 65.3 82.0 / 92.2 90.5 / 92.2 90.2 / 37.1 34.5 / 62.3 23.3 / 77.4 - / 68.0 2.5
StreamingT2V (Henschel et al., 2024) 361 / 330 336 / 331 352 / 245 348 / 258 170 / 150 105 / 98 1672 / 1412 17
90.2 / 84.2 84.0 / 93.4 88.0 / 93.4 87.0 / 57.2 42.5 / 74.7 26.3 / 87.6 - / 84.5 4.25
VideoCrafter2 (Chen et al., 2024a) 382 / 337 324 / 311 356 / 88 365 / 150 98 / 67 64 / 53 1589 / 1006 10
95.5 / 84.8 81.0 / 84.3 89.0 / 84.3 91.2 / 32.1 24.5 / 61.2 16.0 / 76.6 - / 63.3 2.5
Open-sora-plan (Lab and etc., 2024) 379 / 337 342 / 342 357 / 228 367 / 224 134 / 85 68 / 61 1647 / 1277 9
94.8 / 87.9 85.5 / 98.7 89.3 / 98.7 91.7 / 58.3 33.5 / 62.7 17.0 / 88.2 - / 77.6 2.25
MagicTime (Yuan et al., 2024) 320 / 296 267 / 266 299 / 184 303 / 182 248 / 183 216 / 189 1653 / 1300 67
80.0 / 91.2 66.8 / 97.7 74.8 / 97.7 75.7 / 58.4 62.0 / 71.4 54.0 / 86.6 - / 78.7 16.8
StableVideo (AI, 2024a) 393 / 297 369 / 334 384 / 109 381 / 154 128 / 69 72 / 52 1727 / 1015 2
98.3 / 73.5 92.3 / 85.4 96.0 / 85.4 95.3 / 33.3 32.0 / 46.9 18.0 / 65.3 - / 58.8 0.5
Sora (OpenAI, 2024b) 400 / 3 391 / 85 394 / 25 396 / 9 260 / 8 141 / 5 1982 / 135 0
100.0 / 0.75 97.7 / 21.7 98.5 / 6.35 99.0 / 2.27 65.0 / 3.08 35.2 / 3.55 - / 6.81 0.0
Show1 (Zhang et al., 2023a) 382 / 294 355 / 326 372 / 168 380 / 35 128 / 42 70 / 26 1687 / 891 12
95.5 / 77.0 88.8 / 91.3 93.0 / 45.2 95.0 / 9.21 32.0 / 32.8 17.5 / 37.1 - / 52.8 3.0
Dreamina (ByteDance, 2024) 390 / 17 387 / 202 387 / 59 387 / 14 251 / 22 150 / 18 1952 / 332 3
97.5 / 4.36 96.8 / 52.2 96.8 / 15.3 96.8 / 3.62 62.8 / 8.76 37.5 / 12.0 - / 17.0 0.75
Ying (AI, 2024b) 399 / 32 368 / 225 389 / 48 395 / 14 180 / 15 88 / 12 1819 / 346 1
99.8 / 8.02 92.0 / 61.1 97.3 / 12.3 98.8 / 3.54 45.0 / 8.33 22.0 / 13.6 - / 19.0 0.25
LTX (HaCohen et al., 2024) 304 / 196 234 / 226 235 / 162 260 / 67 113 / 59 52 / 35 1198 / 745 89
76.0 / 64.5 58.5 / 96.6 58.8 / 68.9 65.0 / 25.8 28.3 / 52.2 13.0 / 67.3 - / 62.2 22.3
Wan (Cloud, 2025) 394 / 3 379 / 111 387 / 18 389 / 3 154 / 3 53 / 1 1756 / 139 1
98.5 / 0.76 94.8 / 29.3 96.8 / 4.65 97.3 / 0.77 38.5 / 1.95 13.3 / 1.89 - / 7.92 0.25
Kling (Technology, 2024) 393 / 4 367 / 104 385 / 24 393 / 5 222 / 5 123 / 3 1883 / 145 0
98.3 / 1.02 91.8 / 28.3 96.3 / 6.23 98.3 / 1.27 55.5 / 2.25 30.8 / 2.44 - / 7.70 0.0
We calculate the proportion of occurrences and distortions for each body part of AGVs for different T2V models, as detailed in Table 6. For body part occurrences, Sora and Kling perform exceptionally well, with both models including at least one body part in every generated video. Wan, Ying, StableVideo, and Dreamina also demonstrate strong performance, with fewer than 1% of their videos missing any body parts. In contrast, MagicTime and LTX perform poorly, with 16.8% and 22.3% of their videos lacking body parts, respectively. These results indicate that these models struggle with maintaining consistency between text prompts and visual content.
In terms of distortions, Sora, Wan, and Kling show excellent results, particularly Sora, which has only 0.75% of videos with noticeable facial distortions. Ying and Dreamina also exhibit good performance with low distortion rates. However, StableVideo, AnimateDiff, Show1, and Gen-2 show moderate results with some distortions. Open-Sora and MagicTime demonstrate significant distortion, with certain body parts exceeding 95%. This suggests these models struggle with maintaining consistency between text descriptions and the generated visuals.
@@
B.4. The Relationship between Semantic Distortion Identification and Visual Quality Scoring
@@
Figure 13. The correlation between the semantic distortion identification and subjective quality.
@@
B.4.1. Body Parts Distortion
Overall, the quality of T2V models across the three evaluated dimensions demonstrates a certain degree of correlation with the distortions in the five body parts. Among these, facial distortion has the most significant impact on human appearance quality. This pattern aligns with the natural human perception process, which prioritizes facial recognition before focusing on the torso and other body parts.
Of the three quality dimensions, HA Quality exhibits the strongest correlation with body part distortions, followed by AC Quality, with OV Quality showing the weakest correlation. This suggests that body part distortions most directly influence perceptions of human appearance quality, with a lesser effect on the continuity of human actions and overall video quality.
Interestingly, for certain models, such as VideoCrafter2, quality scores show a negative correlation with distortions in the hands and feet. This implies that viewers’ attention is drawn more to prominent areas like the face and other body parts, making distortions in less visually critical body parts less influential.
B.4.2. Body Parts Presence
@@
Interestingly, the presence of legs and feet often correlates negatively with all three quality dimensions, suggesting that their inclusion tends to degrade overall video quality. This underscores the current limitations of T2V models in accurately generating human legs and feet, as distortions in these areas significantly impact viewers’ subjective perception of video quality across three dimensions.
Appendix C Experiment Settings
C.1. Evaluation Metrics
@@
(13) $PLCC=\frac{\sum_{n=1}^{N}(y_{n}-\bar{y})(\hat{y}_{n}-\bar{\hat{y}})}{\sqrt{\sum_{n=1}^{N}(y_{n}-\bar{y})^{2}\sum_{n=1}^{N}(\hat{y}_{n}-\bar{\hat{y}})^{2}}},$
@@
C.2. Details of Evaluation for Distortion Identification
In this study, we obtain outputs from Multi-Modal Large Language Models (MLLMs) through API calls or local deployment. We begin by extracting up to $32$ frames from each video, or all frames if fewer than $32$ are available. These frames, together with a prompt designed for the Distortion Identification task, are then fed into the MLLM. Finally, the raw outputs from the model undergo a post-processing procedure to produce the final results. The detailed prompts are listed below.
For body part occurrences:
# Suppose you are an expert in video content analysis. Please carefully examine and analyze these video frames, and determine whether a person’s [body part] appears in this video. Answer with ”Yes” or ”No”. #.
For body part distortions:
# Suppose you are an expert in video content analysis. Sequentially assemble the provided frames to create a complete video. Then meticulously examine and analyze the video, and determine whether there is any incompleteness, unrealistic appearance, or discontinuous movements of the person’s [body part] in this video. Answer with ”Yes” or ”No”. #.
C.3. Compared Quality Metrics
C.3.1. Image Quality Assessment Methods
@@
NIQE (Mittal et al., 2012b) is a no-reference image quality assessment (IQA) model that evaluates the quality of images based on natural scene statistics (NSS) without relying on human-rated training data. For the Human-AGVQA dataset, we applied NIQE to assess image quality using selected video frames.
@@
BRISQUE (Mittal et al., 2012a) leverages directional contrast features extracted from the Curvelet domain to assess the quality of contrast-distorted images. For the Human-AGVQA dataset, we utilized this approach to evaluate image quality based on its directional contrast features.
@@
HyperIQA (Su et al., 2020) introduces a self-adaptive hyper network for blind image quality assessment (BIQA) in the wild, addressing diverse content and distortion types in real-world images. For the Human-AGVQA dataset, we applied HyperIQA to assess image quality by adaptively learning perception rules and predicting quality scores.
@@
UNIQUE (Zhang et al., 2021) is a unified uncertainty-aware BIQA model designed to handle synthetic and realistic distortions. The model predicts both quality scores and associated uncertainties, enabling a probabilistic interpretation of image quality. For the Human-AGVQA dataset, we used UNIQUE to assess video frame quality based on its learning-to-rank framework.
@@
MUSIQ (Ke et al., 2021) introduces a multi-scale image quality Transformer designed to handle full-resolution images with varying sizes and aspect ratios. The model uses hash-based 2D spatial embedding and scale embedding to capture positional and scale information effectively. For the Human-AGVQA dataset, we applied MUSIQ to evaluate video frame quality by leveraging its multi-scale feature extraction capability.
@@
StairIQA (Sun et al., 2023) introduces a staircase network that integrates features from multiple convolutional layers of a CNN to capture better both low-level visual details and high-level semantic information for in-the-wild image quality assessment. For the Human-AGVQA dataset, we used the official implementation and adopted this approach to evaluate image quality without modifications.
@@
CLIP-IQA (Wang et al., 2023a) is a visual perception assessment model that utilizes the pre-trained CLIP framework to evaluate both the quality and abstract attributes of images. It introduces an antonym prompt pairing strategy (e.g., Good photo vs. Bad photo) to mitigate linguistic ambiguity and employs cosine similarity to predict image quality scores. For the Human-AGVQA dataset, we used the official implementation, applying CLIP-IQA to assess video frame quality.
@@
LIQE (Zhang et al., 2023b) is a multitask learning framework for blind image quality assessment (BIQA) that leverages vision-language correspondence. LIQE uses CLIP for feature embeddings and optimizes tasks jointly by marginalizing probabilities over a textual template that describes scene, distortion, and quality. For the Human-AGVQA dataset, we used the official implementation, applying the quality assessment textual template to evaluate the video frame quality.
@@
MA-AGIQA (Wang et al., 2024b) introduces a framework for AI-generated image quality assessment by combining large multi-modality models (LMMs) with traditional deep neural networks (DNNs). For the Human-AGVQA dataset, we used the official implementation, applying MA-AGIQA to evaluate the video frame quality without modifications.
C.3.2. Video Quality Assessment Methods
@@
TLVQM (Korhonen, 2019) is a two-level video quality model that first computes low-complexity features across the entire video and then selects representative frames based on these features to extract high-complexity features.
@@
RAPIQUE (Tu et al., 2021b) is a video quality assessment model that combines spatial-temporal scene statistics with high-level features extracted using deep CNNs. For the Human-AGVQA dataset, we used the official implementation, applying sparse frame sampling and integrating spatial and temporal features to evaluate video quality.
@@
VIDEVAL (Tu et al., 2021a) applies a feature selection strategy based on efficient blind VQA models. Using the official open-source code, we converted Human-AGVQA videos from RGB to YUV420 format for feature extraction.
@@
Patch-VQ (Ying et al., 2021) is a video quality model that extracts spatial and temporal features from video frames and patches to analyze the relationship between local distortions and overall video quality. For the Human-AGVQA dataset, we used Patch-VQ’s official implementation, dividing each video into space-time patches for feature extraction and quality prediction.
@@
SimpleVOA (Sun et al., 2022) uses an end-to-end network to extract spatial features directly from video frames and motion features to capture temporal distortions. A pre-trained SlowFast model processes 8 uniformly sampled frames rescaled to a height of 520. We used the official SimpleVOA model and finetuned it on Human-AGVQA.
@@
FAST-VQA (Wu et al., 2022) introduces a grid mini-patch sampling (GMS) strategy to balance local quality assessment with global context while reducing computational costs for high-resolution videos. We used the official FAST-VQA-B model and finetuned it on Human-AGVQA.
@@
DOVER (Wu et al., 2023b) is a video quality evaluator that assesses quality from technical and aesthetic perspectives. We used the official code without modifications.
@@
T2VQA (Kou et al., 2024) is a transformer-based model designed to evaluate text-to-video quality, focusing on text-video alignment and video fidelity. It uses BLIP and Swin-T for feature extraction, followed by feature fusion through a cross-attention module and quality regression using a large language model. On the Human-AGVQA dataset, we used the official T2VQA model and finetuned it on Human-AGVQA.
Table 7. Details of the model structure for the proposed model.
@@
UGVQ (Zhang et al., 2024) is a unified framework for video quality assessment, focusing on spatial quality, temporal coherence, and text-to-video alignment. It utilizes features from CLIP for text and visual representations and SlowFast for motion representation. For the Human-AGVQA dataset, we employed UGVQ’s feature extraction and fusion modules to evaluate video quality comprehensively across these three dimensions.
EvalCrafter (Liu et al., 2024a) is an evaluation framework for text-to-video (T2V) generative models, focusing on video quality, text-video alignment, motion quality, and temporal consistency. It uses a benchmark of 700 prompts and 17 evaluation metrics to analyze T2V model performance. For overall video quality, we directly used implementation code in EvalCrafter on the Human-AGVQA dataset without specific changes.
Q-Align (Wu et al., 2023c) introduces a methodology for teaching large multi-modality models (LMMs) to assess video quality using discrete text-defined levels such as excellent, good, and poor. It converts mean opinion scores (MOS) into these rating levels for training and infers scores by weighting the predicted probabilities of the levels. For the Human-AGVQA dataset, we used the official code for inference without modifications.
C.3.3. Action Quality Assessment Methods
USDL (Tang et al., 2020) is an uncertainty-aware score distribution learning approach for action quality assessment (AQA), treating actions as instances associated with score distributions. For the Human-AGVQA dataset, since all videos contain more than 16 frames, each video is evenly divided into ten segments. Feature extraction is performed using an I3D backbone pre-trained on Kinetics. The final scores are normalized linearly, scaling the raw scores based on their minimum and maximum values to a range of [0, 100]. These normalized scores are then used to construct Gaussian distributions with the normalized value as the mean.
ACTION-NET (Zeng et al., 2020) is a hybrid attention network tailored for action quality assessment (AQA) in long videos. It combines dynamic video information with static postures of action subjects in selected frames. For the dynamic stream, frames are sampled at a rate of 4 frames per second. For the static stream, the first, middle, and last frames are sampled, and the action subjects are cropped using the original detection algorithm.
Table 8. Details of the hyper-parameters for the model training.